1. which sets are a basis for the null space of [(1,1,-1,1),(2,1,1,4),(1,0,0,1)].
  2. let T:R^3 to R^3 be defined as T(x,y,z)={x+y, x-y,x+2z). then the basis of range T is...?
  3. let R^3 to R^3 defined as T(x,y,z)=(2x,x+y,x-z). then the adjoint operator T*(u,v,w) is (i) (2u+v+w,v,-w), (ii) (2u,v+w,u-w), (iii) (u,v,-w)
1
Expert's answer
2022-01-03T11:47:07-0500

1.

"A=\begin{pmatrix}
 1 & 1&-1&1 \\
 2 & 1&1&4\\
 1 & 0&0&1\\

\end{pmatrix}"


for null space: "Ax=0"

so, we have:

"x_1=-x_4"

"x_2=x_3"

"-2x_1+2x_2=0\implies x_1=x_2"

solution:

"x=\begin{pmatrix}
 x_1 \\
 x_2\\
x_3\\
x_4
\end{pmatrix}=\begin{pmatrix}
 x_1 \\
 x_1\\
x_1\\
-x_1
\end{pmatrix}=x_1\begin{pmatrix}
 1 \\
 1\\
1\\
-1
\end{pmatrix}"


then basis for the null space:

"\begin{pmatrix}
 1 \\
 1\\
1\\
-1
\end{pmatrix}"


2.

"\begin{pmatrix}
 1 &1&0 \\
 1 &-1&0\\
1 &0&2
\end{pmatrix}\begin{pmatrix}
 x \\
 y\\
z
\end{pmatrix}=\begin{pmatrix}
 x+y \\
 x-y\\
x+2z
\end{pmatrix}"


"\begin{pmatrix}
 1 &1&0 \\
 1 &-1&0\\
1 &0&2
\end{pmatrix}\to \begin{pmatrix}
 1 &1&0 \\
 2 &0&0\\
1 &0&2
\end{pmatrix}"


basis of range T:

"\begin{pmatrix}
 1 \\
 2\\
1
\end{pmatrix},\begin{pmatrix}
 1 \\
 0\\
0
\end{pmatrix},\begin{pmatrix}
 0 \\
 0\\
2
\end{pmatrix}"


3.

i)

Transforming matrix "T" ;

"\begin{pmatrix}
 2x \\
 x+y \\
x-z
\end{pmatrix}" ="x\begin{pmatrix}
 2\\
1\\ 
1
\end{pmatrix}" "+y\begin{pmatrix}
 0 \\
 1\\ 
0
\end{pmatrix}" +"z\begin{pmatrix}
 0 \\
 0\\
-1
\end{pmatrix}"


"T=\begin{pmatrix}
 2& 0&0\\
 1&1 & 0\\
1&0&-1
\end{pmatrix}"


"T^*=T^T=\begin{pmatrix}
 2&1 & 1 \\
 0&1& 0\\
0&0&-1
\end{pmatrix}"


"\begin{pmatrix}
 2&1& 1\\
 0&1 & 0\\
0&0&-1
\end{pmatrix}\begin{pmatrix}
 u \\
 v\\
w
\end{pmatrix}=\begin{pmatrix}
 2u+v+w \\
 v \\
-w
\end{pmatrix}" 


Therefore the adjoint operator "T^*(u,v,w)"

"=(2u+v+w,v,-w)"